Solve All pleasePart A How far must you stretch a spring with ke - 1300 N/m to store 200 J of energy? Express your answer using two significant figures. Submit My Answers Give Up
Answer #1

Here,

part A) let the elongation is x

energy in spring = 0.50 * k * x^2

200 = 0.50 * 1300 * x^2

x = 0.554 m

the elongation is 0.554 m

======================================

part A)

change in thermal energy = decrease in mechanical energy

change in thermal energy = 25 * 9.8 * 3.6 - 0.50 * 25 * 1.7^2

change in thermal energy = 845.9 J

the change in thermal energy is 845.9 J